2. Let X have mean 2 and variance 9. (a) What is the mean and variance of 5-3X? (b) What is the mean and variance of (X - 2)/3? (c) What is the mean and variance of (5X - 10)/2?
Answer #1

X have mean 2 and variance 9

a)

E(5-3x) = 5 - 3 *E(x) = 5-3*2=-1

V(5-3x) = V(-3x)

=9*V(x)

=9*9

=81

b)

T-21

2 1

c)

5r 10 1 EC

5x 10 1 V ( 25 )--V(5x-10)--* 9 = 56.25